OpenStudy (anonymous):

2^3 • 2^4 is equal to... a) seven factors of two b) two factors of seven c) 12 factors of two d) two times 12

OpenStudy (raden):

2^3 * 2^4 = 2^(3+4) = 2^7 how do say it ?

OpenStudy (raden):

* do u